# thanks: Corresponding author

### v1

@@ II.2 Population models of SMBHBs
The differential comoving number density of SMBHBs can be described as ${\rm{d}}^{3}{\Phi_{\mathrm{BHB}}}/{\mathrm{d}}M_{\mathrm{BH},1}\mathrm{d}z\mathrm{d}q$, where $M_{\rm BH,1}$ is the mass of the larger SMBHB, and $q$ is the mass ratio defined as $q=M_{\rm BHB}/M_{\rm BH,1}-1$ with $M_{\mathrm{BHB}}$ being the total binary mass. $q$ is assumed to follow a log-normal distribution, $p(q)$, centered at ${\rm log_{10}}q=0$ with a standard deviation of $\sigma_{{\rm log_{10}}q}=0.5~{}$dex, within the range [0.25, 1.0], similar to the “hiq” model described in Sesana et al. Sesana et al. (2018). While the galactic major merger models can effectively estimate the distribution of SMBHBs, these models are based on the assumptions about the evolution processes of the galaxies and black holes, the galaxy merger rates, and the relationship between the galaxy mass and the black hole mass, etc. However, these assumptions vary greatly among different major merger models, leading to significantly different SMBHB number densities in simulations. It should be noted that the SGWB characteristic amplitude predicted by the M17 model was lower than that of NANOGrav 12.5-yr data. This issue must be addressed if the SGWB signal is considered to be contributed by SMBHBs. We note that the C21 model Casey-Clyde et al. (2022) can self-consistently incorporate the SGWB characteristic amplitude as a parameter to infer the distribution of SMBHBs, and the number of nearby SMBHBs inferred from the NANOGrav 12.5-yr data is more than five times the number estimated by the M17 major merger model. Therefore, we choose the C21 model to estimate the distribution of SMBHBs and compare it with the M17 major merger model Yan et al. (2020); Wang et al. (2022d).
II.2.1 Major merger model
@@
We convert $M_{K}$ to the stellar masses for early-type galaxies using the empirical equation, ${\rm log}_{10}M_{\star}=10.58-0.44(M_{K}+23)$ Cappellari (2013). The masses of SMBHBs in these galaxies are determined via the $M_{\rm BHB}$-$M_{\rm bulge}$ relationship from Ref. McConnell and Ma (2013), where the bulge mass, $M_{\rm bulge}$, is approximated to the stellar mass $M_{\star}$. The mass ratio $q$ of the binary is drawn from a log-normal distribution, $p(q)$, centered at ${\rm log_{10}}q$ = 0 and with the standard deviation of $\sigma_{{\rm log_{10}}q}$ = 0.5 dex, within the range [0.25, 1.0]. Then, the probability of the $j$-th galaxy hosting an SMBHB in the PTA band, $p_{j}$, is given by
$p_{j}({M_{\star}},{z})=\frac{t_{{\rm c},j}}{T_{{\rm life}}}\int_{\mu_{\star}\geq 0.25}\mathrm{~{}d}\mu_{\star}\frac{\mathrm{d}N}{\mathrm{~{}d}t}\left({M_{\star}},\mu_{\star},{z}^{\prime}\right)T_{{\rm life}},$ (2)
@@
where $t_{{\rm c},j}=(5/256)(\pi f_{\rm low})^{-8/3}(GM_{\rm c}/c^{3})^{-5/3}$ is the time to coalescence of the SMBHB in the $j$-th galaxy, with $f_{\rm low}=1$ nHz being the lower limit of the PTA band. $T_{\rm life}=T_{\rm df}+T_{\rm sh}$ is the lifetime of the SMBHB, where $T_{\rm df}$ and $T_{\rm sh}$ are the dynamical friction timescale Binney and Tremaine (2008) and the stellar hardening timescale Sesana and Khan (2015), respectively. $\mathrm{d}N/\mathrm{d}t\left(M_{\star},\mu_{\star},z^{\prime}\right)$ represents the galaxy merger rate derived from the Illustris cosmological simulation project Genel et al. (2014); Rodriguez-Gomez et al. (2015), where $M_{\star}$ denotes the stellar mass of the galaxies, $\mu_{\star}$ denotes the progenitor stellar mass ratio, and $z^{\prime}$ denotes the redshift at which the galaxies merge, calculated at $T_{z}$ with Planck cosmological parameters Ade et al. (2016).
@@
II.2.2 Quasar-based model
The characteristic amplitude $h_{\rm c}$ of the GW signal generated by the population of merging SMBHBs can be expressed as Phinney (2001); Sesana et al. (2009); Sesana (2013)
$h_{\rm c}^{2}(f)=\frac{4}{\pi f^{2}}\iiint$ ${\rm{d}}M_{\mathrm{BH},1}{\rm{d}}z{\rm{d}}q$ (3)
×$\times\frac{{\rm{d}}^{3}\Phi_{\mathrm{BHB}}}{{\rm{d}}M_{\mathrm{BH},1}{\rm{d}}z{\rm{d}}q}\frac{1}{1+z}\frac{{\rm{d}}E_{\mathrm{gw}}}{{\rm{d}}\ln f_{\rm{r}}},$
where $\rm{d}^{3}\Phi_{\mathrm{BHB}}/d\textit{M}_{\mathrm{BH},1}\,d\textit{z}\,d\textit{q}$ represents the differential comoving number density of SMBHBs, ${\rm d}E_{\mathrm{gw}}/{\rm d}\ln f_{\rm r}$ is the GW energy emitted per unit logarithmic frequency change, and $f_{\rm r}=f(1+z)$ is the GW frequency in the rest frame of the source. For a population of SMBHBs in circular orbits, the GW energy per unit logarithmic frequency change is Thorne (1987); Phinney (2001)
$\frac{{\rm d}E_{\mathrm{gw}}}{{\rm d}\ln f_{\rm{r}}}$ $=\frac{\left(\pi f_{\rm r}\right)^{2/3}}{3}{M_{\rm c}}^{5/3},$ (4)
@@ II.2.2 Quasar-based model
$\frac{\rm{d}^{3}\Phi_{\mathrm{BHB}}}{{\rm{d}}M_{\mathrm{BH},1}{\rm{d}}z{\rm{d}}q}\propto\frac{\rm{d}^{2}\Phi_{\mathrm{Q}}}{{\rm{d}}M_{\mathrm{BH}}{\rm{d}}z}p(q).$ (5)
That is to say, the differential comoving number density of SMBHBs is proportional to the differential comoving number density of quasars, ${{\rm{d}}^{2}\Phi_{\mathrm{Q}}}/{{\rm{d}}M_{\mathrm{BH}}{\rm d}z}$, and an assumed mass ratio distribution $p(q)$. The number density of quasars is derived from the quasar luminosity function and Eddington ratio distribution of the lifetime of the quasar from the galaxy merger simulations in Ref. Hopkins et al. (2006). To estimate the number density of quasars, we convert the differential quasar triggering rate Hopkins et al. (2007), ${\mathrm{d}^{2}\Phi_{\mathrm{Q}}}/{\mathrm{d}M_{\mathrm{BH}}{\mathrm{d}}t_{\rm r}}$, into a redshift distribution using the $\Lambda$CDM model with WMAP9 data,
$\frac{\mathrm{d}^{2}\Phi_{\mathrm{Q}}}{\mathrm{d}M_{\mathrm{BH}}\mathrm{d}z}=\frac{\mathrm{d}^{2}\Phi_{\mathrm{Q}}}{\mathrm{d}M_{\mathrm{BH}}\mathrm{d}t_{\rm r}}\frac{\mathrm{d}t_{\rm r}}{\mathrm{d}z}.$ (6)
According to Eqs. (5) and (6), we can renormalize the SMBHB population with both the SGWB characteristic amplitude and the quasar number density distribution,
$\frac{\mathrm{d}^{3}\Phi_{\mathrm{BHB}}}{\mathrm{d}M_{\mathrm{BH},1}\mathrm{d}z\mathrm{d}q}=\frac{\Phi_{\mathrm{BHB},0}}{\Phi_{\mathrm{Q},0}}\frac{\mathrm{d}^{2}\Phi_{\mathrm{Q}}}{\mathrm{d}M_{\mathrm{BH}}\mathrm{d}t_{\rm r}}\frac{\mathrm{d}t_{\rm r}}{\mathrm{d}z}\frac{p(q)}{\int_{q\geq 0.25}\mathrm{d}qp(q)},$ (7)
@@
$\Phi_{\mathrm{BHB},0}=$ $\left.\iint_{\begin{subarray}{c}M_{\mathrm{BH},1}\geq M_{\mathrm{BH},\mathrm{det}},\\ q\geq 0.25\end{subarray}}\mathrm{d}M_{\mathrm{BH}}\mathrm{d}q\frac{\mathrm{d}^{3}\Phi_{\mathrm{BHB}}}{\mathrm{d}M_{\mathrm{BH},1}\mathrm{d}z\mathrm{d}q}\right|_{z=0},$ (8)
$\Phi_{\mathrm{Q},0}=\left.\int_{M_{\mathrm{BH}}\geq M_{\mathrm{BH},\mathrm{det}}}\mathrm{d}M_{\mathrm{BH}}\frac{\mathrm{d}^{2}\Phi_{\mathrm{Q}}}{\mathrm{d}M_{\mathrm{BH}}\mathrm{d}t_{\rm r}}\frac{\mathrm{d}t_{\rm r}}{\mathrm{d}z}\right|_{z=0}.$ (9)
The best-fit value of the SGWB characteristic amplitude from the NANOGrav 15-yr data set is $h_{\rm c}=2.4\times 10^{-15}$ Agazie et al. (2023a). Assuming that this SGWB signal is contributed by numerous SMBHB inspirals, according to our calculations using Eqs. (3)–(9), there are approximately 688 SMBHBs in the redshift range (0, 0.1). Considering the completeness of the 2MASS catalog, we use Eq. (1) to calculate the total completeness fraction of the 688 SMBHBs. This completeness fraction actually refers to the proportion of the host galaxies included in the 2MASS catalog relative to the total host galaxies. We find that it is approximately equal to 490. Therefore, we randomly select 490 galaxies based on their $P_{\rm com}$ values from the 2MASS catalog as SMBHBs’ host galaxies. We assume that the host galaxies of the other 198 SMBHBs are not included in the 2MASS catalog and ignore their contributions to the subsequent calculations for dark sirens.
@@
II.3 GW detection and Fisher analysis
@@
$s(t,\hat{\Omega})=F^{+}(\hat{\Omega})\Delta A_{+}(t)+F^{\times}(\hat{\Omega})\Delta A_{\times}(t).$ (10)
Here, $F^{+}(\hat{\Omega})$ and $F^{\times}(\hat{\Omega})$ are the antenna pattern functions Yan et al. (2020), which can be expressed as
@@
where ($\alpha$, $\delta$) and ($\alpha_{\rm p}$, $\delta_{\rm p}$) represent the right ascension and declination of the GW source and the pulsar, respectively. $\theta$ represents the opening angle between the GW source and the pulsar,
$\cos\theta=\cos\delta\cos\delta_{\mathrm{p}}\cos\left(\alpha-\alpha_{\mathrm{p}}\right)+\sin\delta\sin\delta_{\mathrm{p}}.$ (12)
In Eq. (10), $\Delta A_{\{+,\times\}}(t)=A_{\{+,\times\}}(t)-A_{\{+,\times\}}\left(t_{\mathrm{p}}\right)$ Ellis et al. (2012) represents the difference between the Earth term $A_{\{+,\times\}}(t)$ and the pulsar term $A_{\{+,\times\}}(t_{\mathrm{p}})$. $t_{\rm{p}}$ is the time when the GW passes the pulsar and can be expressed as $t_{\rm{p}}=t-d_{\rm{p}}(1-{\rm{cos}}\theta)/c$, with $d_{\rm{p}}$ representing the pulsar distance. For evolving SMBHBs in circular orbits, we have
@@
$h(t)=2\frac{\left(GM_{\mathrm{c}}\right)^{5/3}}{c^{4}}\frac{(\pi f(t))^{2/3}}{d_{\mathrm{L}}},$ (15)
where the frequency of GW signals is calculated by the following formula Peters (1964),
$f(t)=\pi^{-1}\left(\frac{G{M}_{\rm c}}{c^{3}}\right)^{-5/8}\left[\frac{256}{5}(t_{\rm c}-t)\right]^{-3/8}.$ (16)
Here, $t_{\rm c}$ is uniformly sampled from [100 yr, 26 Myr]. This range represents the time to coalescence of a SMBHB with $M_{\rm BH,1}=10^{9}M_{\odot}$ and $q=1$ from 1 nHz to 100 nHz. $h_{0}$ and $f_{0}$ represent the values of $h(t)$ and $f(t)$ when $t=0$, respectively. Fig. 2 shows the distribution of mock SMBHBs on the $f_{0}$-$h_{0}$, $f_{0}$-$M_{\rm BHB}$, $z$-$M_{\rm BHB}$ planes. The inclination angle, the polarization angle, and the initial phase are randomly sampled in the ranges $\cos\iota\in[-1,1]$, $\psi\in[0,2\pi]$, and $\phi_{0}\in[0,2\pi]$, respectively. The GW phase, $\phi(t)$, can be expressed as
$\phi(t)=\phi_{0}+\frac{1}{16}\left(\frac{GM_{\mathrm{c}}}{c^{3}}\right)^{-5/3}\left\{\left(\pi f_{0}\right)^{-5/3}-[\pi f(t)]^{-5/3}\right\}.$ (17)
@@
$\rho^{2}=\sum_{j=1}^{N_{\mathrm{p}}}\sum_{i=1}^{N}\left[\frac{s_{j}\left(t_{i}\right)}{\sigma_{t,j}}\right]^{2},$ (18)
where $N$ is the total number of the data points of timing residuals for each pulsar, $N_{\rm p}$ is the number of pulsars, $s_{j}(t_{i})$ is the timing residual induced by the GW signal in the $j$-th pulsar at time $t_{i}$, and $\sigma_{t,j}$ is the RMS of the white noise of the $j$-th pulsar. The SNR threshold is set to 10 in the simulation.
It should be noted that in this work, we did not consider red noise, nor did we take an unresolved SGWB as red noise into account. We assume that the pulsar noise is primarily dominated by white noise and that individual GW sources can be disentangled from the SGWB through specialized methodologies.
We utilize the Fisher information matrix for parameter estimations. For a PTA comprising $N_{\rm p}$ independent pulsars, the Fisher information matrix can be expressed as
$\Gamma_{ab}=\sum_{j=1}^{N_{\rm p}}\sum_{i=1}^{N}\frac{\partial s\left(t_{i}\right)}{\sigma_{t,j}\partial\theta_{a}}\frac{\partial s\left(t_{i}\right)}{\sigma_{t,j}\partial\theta_{b}},$ (19)
@@
$\Delta\theta_{a}=\left(\Gamma^{-1}\right)_{aa}^{1/2}.$ (20)
@@
$\Delta\Omega=2\pi\sqrt{(\cos\delta\Delta\delta\Delta\alpha)^{2}-\left(\cos\delta\frac{\left(\Gamma^{-1}\right)_{\delta\alpha}}{\Delta\delta\Delta\alpha}\right)^{2}}.$ (21)
For the parameter $d_{\rm L}$, we first estimate its instrumental error by Eq. (20). Besides, we also consider the error arising from weak lensing, which is expressed by the fitting formula from Refs. Hirata et al. (2010); Tamanini et al. (2016),
$\Delta{d_{\mathrm{L}}^{\text{lens}}}(z)$ $=d_{\mathrm{L}}(z)\times 0.066\left[\frac{1-(1+z)^{-0.25}}{0.25}\right]^{1.8}.$ (22)

### v2

@@ 2.2 Population models of SMBHBs
The differential comoving number density of SMBHBs can be described as ${\rm{d}}^{3}{\Phi_{\mathrm{BHB}}}/{\mathrm{d}}M_{\mathrm{BH},1}\mathrm{d}z\mathrm{d}q$, where $M_{\rm BH,1}$ is the mass of the larger SMBHB, and $q$ is the mass ratio defined as $q=M_{\rm BHB}/M_{\rm BH,1}-1$ with $M_{\mathrm{BHB}}$ being the total binary mass. $q$ is assumed to follow a log-normal distribution, $p(q)$, centered at $\log_{10}q=0$ with a standard deviation of $\sigma_{\log_{10}q}=0.5~{}$dex, within the range [0.25, 1.0], similar to the “hiq” model described in Sesana et al. [156]. While the galactic major merger models can effectively estimate the distribution of SMBHBs, these models are based on the assumptions about the evolution processes of the galaxies and black holes, the galaxy merger rates, and the relationship between the galaxy mass and the black hole mass, etc. However, these assumptions vary greatly among different major merger models, leading to significantly different SMBHB number densities in simulations. It should be noted that the SGWB characteristic amplitude predicted by the M17 model was lower than that of NANOGrav 12.5-yr data. This issue must be addressed if the SGWB signal is considered to be contributed by SMBHBs. We note that the C21 model [151] can self-consistently incorporate the SGWB characteristic amplitude as a parameter to infer the distribution of SMBHBs, and the number of nearby SMBHBs inferred from the NANOGrav 12.5-yr data is more than five times the number estimated by the M17 major merger model. Therefore, we choose the C21 model to estimate the distribution of SMBHBs and compare it with the M17 major merger model [148, 149].
2.2.1 Major merger model
@@
We convert $M_{K}$ to the stellar masses for early-type galaxies using the empirical equation, ${\rm log}_{10}M_{\star}=10.58-0.44(M_{K}+23)$ [157]. The masses of SMBHBs in these galaxies are determined via the $M_{\rm BHB}$-$M_{\rm bulge}$ relationship from McConnell et al. [158], where the bulge mass, $M_{\rm bulge}$, is approximated to the stellar mass $M_{\star}$. The mass ratio $q$ of the binary is drawn from a log-normal distribution, $p(q)$, centered at ${\rm log_{10}}q$ = 0 and with the standard deviation of $\sigma_{{\rm log_{10}}q}$ = 0.5 dex, within the range [0.25, 1.0]. Then, the probability of the $j$-th galaxy hosting an SMBHB in the PTA band, $p_{j}$, is given by
$p_{j}({M_{\star}},{z})=\frac{t_{{\rm c},j}}{T_{{\rm life}}}\int_{\mu_{\star}\geq 0.25}\mathrm{~{}d}\mu_{\star}\frac{\mathrm{d}N}{\mathrm{~{}d}t}\left({M_{\star}},\mu_{\star},{z}^{\prime}\right)T_{{\rm life}},$ (2.2)
@@
where $t_{{\rm c},j}=(5/256)(\pi f_{\rm low})^{-8/3}(GM_{\rm c}/c^{3})^{-5/3}$ is the time to coalescence of the SMBHB in the $j$-th galaxy, with $f_{\rm low}=1$ nHz being the lower limit of the PTA band. $T_{\rm life}=T_{\rm df}+T_{\rm sh}$ is the lifetime of the SMBHB, where $T_{\rm df}$ and $T_{\rm sh}$ are the dynamical friction timescale [159] and the stellar hardening timescale [160], respectively. $\mathrm{d}N/\mathrm{d}t\left(M_{\star},\mu_{\star},z^{\prime}\right)$ represents the galaxy merger rate derived from the Illustris cosmological simulation project [161, 162], where $M_{\star}$ denotes the stellar mass of the galaxies, $\mu_{\star}$ denotes the progenitor stellar mass ratio, and $z^{\prime}$ denotes the redshift at which the galaxies merge, calculated at $T_{z}$ with Planck cosmological parameters [163].
@@
2.2.2 Quasar-based model
The characteristic amplitude $h_{\rm c}$ of the GW signal generated by the population of merging SMBHBs can be expressed as [164, 138, 165]
$h_{\rm c}^{2}(f)=\frac{4}{\pi f^{2}}\iiint$ ${\rm{d}}M_{\mathrm{BH},1}{\rm{d}}z{\rm{d}}q\frac{{\rm{d}}^{3}\Phi_{\mathrm{BHB}}}{{\rm{d}}M_{\mathrm{BH},1}{\rm{d}}z{\rm{d}}q}\frac{1}{1+z}\frac{{\rm{d}}E_{\mathrm{gw}}}{{\rm{d}}\ln f_{\rm{r}}},$ (2.3)
where $\rm{d}^{3}\Phi_{\mathrm{BHB}}/d\textit{M}_{\mathrm{BH},1}\,d\textit{z}\,d\textit{q}$ represents the differential comoving number density of SMBHBs, ${\rm d}E_{\mathrm{gw}}/{\rm d}\ln f_{\rm r}$ is the GW energy emitted per unit logarithmic frequency change, and $f_{\rm r}=f(1+z)$ is the GW frequency in the rest frame of the source. For a population of SMBHBs in circular orbits, the GW energy per unit logarithmic frequency change is [166, 164]
$\frac{{\rm d}E_{\mathrm{gw}}}{{\rm d}\ln f_{\rm{r}}}$ $=\frac{\left(\pi f_{\rm r}\right)^{2/3}}{3}{M_{\rm c}}^{5/3},$ (2.4)
@@ 2.2.2 Quasar-based model
$\frac{\rm{d}^{3}\Phi_{\mathrm{BHB}}}{{\rm{d}}M_{\mathrm{BH},1}{\rm{d}}z{\rm{d}}q}\propto\frac{\rm{d}^{2}\Phi_{\mathrm{Q}}}{{\rm{d}}M_{\mathrm{BH}}{\rm{d}}z}p(q).$ (2.5)
That is to say, the differential comoving number density of SMBHBs is proportional to the differential comoving number density of quasars, ${{\rm{d}}^{2}\Phi_{\mathrm{Q}}}/{{\rm{d}}M_{\mathrm{BH}}{\rm d}z}$, and an assumed mass ratio distribution $p(q)$. The number density of quasars is derived from the quasar luminosity function and Eddington ratio distribution of the lifetime of the quasar from the galaxy merger simulations in Hopkins et al. [167]. To estimate the number density of quasars, we convert the differential quasar triggering rate [168], ${\mathrm{d}^{2}\Phi_{\mathrm{Q}}}/{\mathrm{d}M_{\mathrm{BH}}{\mathrm{d}}t_{\rm r}}$, into a redshift distribution using the $\Lambda$CDM model with WMAP9 data,
$\frac{\mathrm{d}^{2}\Phi_{\mathrm{Q}}}{\mathrm{d}M_{\mathrm{BH}}\mathrm{d}z}=\frac{\mathrm{d}^{2}\Phi_{\mathrm{Q}}}{\mathrm{d}M_{\mathrm{BH}}\mathrm{d}t_{\rm r}}\frac{\mathrm{d}t_{\rm r}}{\mathrm{d}z}.$ (2.6)
According to Eqs. (2.5) and (2.6), we can renormalize the SMBHB population with both the SGWB characteristic amplitude and the quasar number density distribution,
$\frac{\mathrm{d}^{3}\Phi_{\mathrm{BHB}}}{\mathrm{d}M_{\mathrm{BH},1}\mathrm{d}z\mathrm{d}q}=\frac{\Phi_{\mathrm{BHB},0}}{\Phi_{\mathrm{Q},0}}\frac{\mathrm{d}^{2}\Phi_{\mathrm{Q}}}{\mathrm{d}M_{\mathrm{BH}}\mathrm{d}t_{\rm r}}\frac{\mathrm{d}t_{\rm r}}{\mathrm{d}z}\frac{p(q)}{\int_{q\geq 0.25}\mathrm{d}qp(q)},$ (2.7)
@@
$\Phi_{\mathrm{BHB},0}=$ $\left.\iint_{\begin{subarray}{c}M_{\mathrm{BH},1}\geq M_{\mathrm{BH},\mathrm{det}},\\ q\geq 0.25\end{subarray}}\mathrm{d}M_{\mathrm{BH}}\mathrm{d}q\frac{\mathrm{d}^{3}\Phi_{\mathrm{BHB}}}{\mathrm{d}M_{\mathrm{BH},1}\mathrm{d}z\mathrm{d}q}\right|_{z=0},$ (2.8)
$\Phi_{\mathrm{Q},0}=\left.\int_{M_{\mathrm{BH}}\geq M_{\mathrm{BH},\mathrm{det}}}\mathrm{d}M_{\mathrm{BH}}\frac{\mathrm{d}^{2}\Phi_{\mathrm{Q}}}{\mathrm{d}M_{\mathrm{BH}}\mathrm{d}t_{\rm r}}\frac{\mathrm{d}t_{\rm r}}{\mathrm{d}z}\right|_{z=0}.$ (2.9)
In this model, the characteristic amplitude of the SGWB, $h_{c}$, is treated as an input parameter. The Bayesian posterior median of the SGWB characteristic amplitude from the NANOGrav 15-yr dataset is $2.4\times 10^{-15}$ [130], assuming a spectral index of $13/3$, as expected for SMBHB inspirals. We substitute $h_{c}=2.4\times 10^{-15}$ into Eq. (2.3) and use Eqs. (2.3)–(2.9) to calculate the number of SMBHBs. We estimate that there are approximately 688 SMBHBs in the redshift range (0, 0.1). It should be added that using different values of $h_{c}$ results in different numbers of SMBHBs but the C21 model is applicable to the observational signals from any PTA project.
The possible origins of the current signal observed by PTAs can be broadly classified into astrophysical and cosmological sources. In this study, we adopt an astrophysical interpretation, using $h_{c}$ as an input to the C21 model under the assumption that the NANOGrav 15-yr signal is attributed to an ensemble of inspiraling SMBHBs. However, the true origin of the PTA signals remains uncertain. While an incoherent superposition of GWs from numerous SMBHBs provides a natural astrophysical explanation, several alternative cosmological scenarios have been proposed. These include primordial GWs generated by cosmic inflation [169], scalar-induced GWs [170, 171], GWs produced during cosmological first-order phase transitions [172, 173], GW radiation from cosmic strings [174, 175] and domain walls [176], among others. Each of these mechanisms could contribute to the GW signal in the nHz frequency range. The current observational data are insufficient to distinguish these potential sources, making future high-precision observations crucial for uncovering the true origin of this signal.
Considering the completeness of the 2MASS catalog, we use Eq. (2.1) to calculate the total completeness fraction of the 688 SMBHBs. This completeness fraction actually refers to the proportion of the host galaxies included in the 2MASS catalog relative to the total host galaxies. We find that it is approximately equal to 490. Therefore, we randomly select 490 galaxies based on their $P_{\rm com}$ values from the 2MASS catalog as SMBHBs’ host galaxies. We assume that the host galaxies of the other 198 SMBHBs are not included in the 2MASS catalog and ignore their contributions to the subsequent calculations for dark sirens.
@@
2.3 GW detection and Fisher analysis
@@
$s(t,\hat{\Omega})=F^{+}(\hat{\Omega})\Delta A_{+}(t)+F^{\times}(\hat{\Omega})\Delta A_{\times}(t).$ (2.10)
Here, $F^{+}(\hat{\Omega})$ and $F^{\times}(\hat{\Omega})$ are the antenna pattern functions [148], which can be expressed as
@@
where ($\alpha$, $\delta$) and ($\alpha_{\rm p}$, $\delta_{\rm p}$) represent the right ascension and declination of the GW source and MSP, respectively. $\theta$ represents the opening angle between the GW source and MSP,
$\cos\theta=\cos\delta\cos\delta_{\mathrm{p}}\cos\left(\alpha-\alpha_{\mathrm{p}}\right)+\sin\delta\sin\delta_{\mathrm{p}}.$ (2.12)
In Eq. (2.10), $\Delta A_{\{+,\times\}}(t)=A_{\{+,\times\}}(t)-A_{\{+,\times\}}\left(t_{\mathrm{p}}\right)$ [143] represents the difference between the Earth term $A_{\{+,\times\}}(t)$ and the pulsar term $A_{\{+,\times\}}(t_{\mathrm{p}})$. $t_{\rm{p}}$ is the time when the GW passes MSP and can be expressed as $t_{\rm{p}}=t-d_{\rm{p}}(1-{\rm{cos}}\theta)/c$, with $d_{\rm{p}}$ representing the pulsar distance. For evolving SMBHBs in circular orbits, we have
@@
$h(t)=2\frac{\left(GM_{\mathrm{c}}\right)^{5/3}}{c^{4}}\frac{(\pi f(t))^{2/3}}{d_{\mathrm{L}}},$ (2.15)
where the frequency of GW signals is calculated by the following formula [177],
$f(t)=\pi^{-1}\left(\frac{G{M}_{\rm c}}{c^{3}}\right)^{-5/8}\left[\frac{256}{5}(t_{\rm c}-t)\right]^{-3/8}.$ (2.16)
Here, $t_{\rm c}$ is uniformly sampled from [100 yr, 26 Myr]. This range represents the time to coalescence of a SMBHB with $M_{\rm BH,1}=10^{9}M_{\odot}$ and $q=1$ from 1 nHz to 100 nHz. $h_{0}$ and $f_{0}$ represent the values of $h(t)$ and $f(t)$ when $t=0$, respectively. Figure 2 shows the distribution of mock SMBHBs on the $f_{0}$-$h_{0}$, $f_{0}$-$M_{\rm BHB}$, $z$-$M_{\rm BHB}$ planes. The inclination angle, the polarization angle, and the initial phase are randomly sampled in the ranges $\cos\iota\in[-1,1]$, $\psi\in[0,2\pi]$, and $\phi_{0}\in[0,2\pi]$, respectively. The GW phase, $\phi(t)$, can be expressed as
$\phi(t)=\phi_{0}+\frac{1}{16}\left(\frac{GM_{\mathrm{c}}}{c^{3}}\right)^{-5/3}\left\{\left(\pi f_{0}\right)^{-5/3}-[\pi f(t)]^{-5/3}\right\}.$ (2.17)
@@
$\rho^{2}=\sum_{j=1}^{N_{\mathrm{p}}}\sum_{i=1}^{N}\left[\frac{s_{j}\left(t_{i}\right)}{\sigma_{t,j}}\right]^{2},$ (2.18)
where $N$ is the total number of the data points of timing residuals for each MSP, $N_{\rm p}$ is the number of MSPs, $s_{j}(t_{i})$ is the timing residual induced by the GW signal in the $j$-th MSP at time $t_{i}$, and $\sigma_{t,j}$ is the RMS of the white noise of the $j$-th MSP. The SNR threshold is set to 10 in the simulation.
As this paper focuses on the impact of SMBHB population models on dark sirens, we consider only white noise as the noise source, without taking into account red noise or SGWBs. According to the analyses in other papers, taking into account red noise and SGWB could reduce SNRs of GW signals [178], and red noise could worsen the measurement precision of cosmological parameters by 3–5 times [149]. Therefore, we can roughly estimate that, red noise and SGWB would also worsen the constraint precision of $H_{0}$ presented in this paper by a few times. A more detailed analysis of red noise and SGWB will be addressed in our future works.
We utilize the Fisher information matrix for parameter estimations. For a PTA comprising $N_{\rm p}$ independent MSPs, the Fisher information matrix can be expressed as
$\Gamma_{ab}=\sum_{j=1}^{N_{\rm p}}\sum_{i=1}^{N}\frac{\partial s\left(t_{i}\right)}{\sigma_{t,j}\partial\theta_{a}}\frac{\partial s\left(t_{i}\right)}{\sigma_{t,j}\partial\theta_{b}},$ (2.19)
@@
For each GW source, the timing residuals depend on $N_{\rm{p}}+8$ system parameters, including eight parameters of the GW source (i.e., $M_{\rm c},\alpha,\delta,\iota,\psi,\phi_{0},f_{0},d_{\rm L}$) and the distances of MSPs $d_{\rm{p},\textit{j}}$ $(j=1,2,\cdots,N_{\rm p})$. With additional EM information priors, the estimation precision of the parameter can be improved by adding to the appropriate diagonal element of the Fisher information matrix, i.e.,
$\Gamma_{ab}\rightarrow\Gamma_{ab}-\left\langle\frac{\partial^{2}\ln P\left(\theta_{i}\right)}{\partial\theta_{a}\partial\theta_{b}}\right\rangle,$ (2.20)
@@
$\Delta\theta_{a}=\left(\Gamma^{-1}\right)_{aa}^{1/2}.$ (2.21)
@@
$\Delta\Omega=2\pi\sqrt{(\cos\delta\Delta\delta\Delta\alpha)^{2}-\left(\cos\delta\frac{\left(\Gamma^{-1}\right)_{\delta\alpha}}{\Delta\delta\Delta\alpha}\right)^{2}}.$ (2.22)
For the parameter $d_{\rm L}$, we first estimate its instrumental error by Eq. (2.21). Besides, we also consider the error arising from weak lensing, which is expressed by the fitting formula [180, 181],
$\Delta{d_{\mathrm{L}}^{\text{lens}}}(z)$ $=d_{\mathrm{L}}(z)\times 0.066\left[\frac{1-(1+z)^{-0.25}}{0.25}\right]^{1.8}.$ (2.23)